# An empirical appraisal of methods for the dynamic prediction of survival with numerous longitudinal predictors

### v1

@@ 2.3.1 Static Cox model
$h_{i}(t)=h_{0}(t)\exp\left\{\alpha x_{i}+\gamma y_{i}(0)\right\},\>i\in\mathcal{I}(\ell),$ (2)
where $h_{i}(t)$ denotes the hazard for subject $i$ at time $t$, $h_{0}(t)$ is a non-parametric baseline hazard, and $\alpha$ and $\gamma$ are vectors of regression coefficients. Model (2) can be estimated using the R package survival (Therneau and Grambsch,, 2000). The predicted conditional survival probability of equation (1) can be estimated by computing
@@
Admittedly, model (2) does not make use of the available longitudinal measurements to update predictions of survival: nevertheless, it is useful to evaluate the predictive performance of such model to quantify how accurate our predictions of survival would be if we were to ignore any information collected after study entrance. By comparing the performance of this static prediction model to that of dynamic methods, which incorporate more and more repeated measurements as the landmark time increases, we can effectively quantify the extent to which adding longitudinal information to a prediction model may improve predictive performance. Although the discrepancy between static and dynamic methods will be typically data-dependent, in general we can expect the difference between these two approaches to increase as the landmark time increases, as the information provided to model (2) becomes more and more outdated, while dynamic methods continue being fed more and more up to date patient information.
@@ 2.3.2 Landmarking
Landmarking (Van Houwelingen,, 2007) is a pragmatic approach to dynamic prediction that summarizes the trajectory of each longitudinal covariate up until the landmark time into a single summary measure per covariate. Two commonly-used summary measures are the last available observation (up to $\ell$) of the longitudinal covariate, an approach referred to as last observation carried forward (LOCF), and the average of all repeated measurements taken up until $\ell$. The summaries of the longitudinal covariates are then included as predictors in a Cox model alongside with the baseline covariates.
The LOCF landmarking method can be formalized as follows. For each subject $i\in\mathcal{I}(\ell)$ and each longitudinal covariate $q\in\{1,...,Q\}$, consider the full trajectory $y_{qi}=(y_{qi1},y_{qi2},...,y_{qim_{i}})$ and denote by $y_{qi}^{\ell}$ the last non-missing value $y_{qik}$ in $y_{qi}$ such that $t_{ik}\leq\ell$ (notice that if there are no missing values in $y_{qi}$ for every $q$, then the measurement times of $y_{qi}^{\ell}$ are the same across $q$, otherwise they may differ). The LOCF approach involves fitting the Cox model
$h_{i}(t)=h_{0}(t)\exp\left\{\alpha x_{i}+\sum_{q=1}^{Q}\gamma_{q}y_{qi}^{\ell}\right\},\>i\in\mathcal{I}(\ell),$ (4)
where $\alpha$ and $\gamma=(\gamma_{1},\gamma_{2},...,\gamma_{Q})$ are vectors of coefficients.
The LOCF landmarking model can be easily estimated using existing software for survival analysis, such as the R package survival (Therneau and Grambsch,, 2000). The predicted survival for subject $i$ can be estimated using equation (3), where now $\hat{\eta}_{i}$ is the estimated linear predictor for subject $i$ from model (4).
@@ 2.3.3 Multivariate Functional Principal Component Analysis Cox model (MFPCCox)
The Multivariate Functional Principal Component Analysis Cox (MFPCCox) model (Li and Luo,, 2019) is a prediction approach that uses Multivariate Functional Principal Component Analysis (MFPCA) to summarize the $Q$ longitudinal covariates into MFPCA scores, which are then used as predictors of survival in a Cox model.
@@
where $\psi_{k}(t)=\left(\psi_{k1}(t),...,\psi_{kQ}(t)\right)$ are $Q$-variate orthonormal eigenfunctions with associated MFPCA scores $\rho_{ki}$. The eigenfunctions are ordered by decreasing percentage of variance explained, and $K$ is chosen in such a way that the first $K$ eigenfunctions explain a certain proportion of the variance of the longitudinal covariates.
To predict survival, MFPCCox uses the estimated MFPCA scores for subject $i$ as summaries of all repeated measurements of the longitudinal covariates, including the MFPCA scores alongside the baseline covariates as predictors in a Cox model:
$h_{i}(t)=h_{0}(t)\exp\left\{\alpha x_{i}+\sum_{k=1}^{K}\gamma_{k}\hat{\rho}_{ki}\right\},\>i\in\mathcal{I}(\ell),$ (6)
where $\hat{\rho}_{ki}$ are the estimated MFPCA scores, and $\alpha$ and $\gamma=(\gamma_{1},\gamma_{2},...,\gamma_{K})$ are vectors of coefficients. MFPCCox can be estimated using the R code provided in Li and Luo, (2019). The predicted survival for subject $i$ can be estimated using equation (3), where $\hat{\eta}_{i}$ is now the estimated linear predictor for subject $i$ from model (6).
@@ 2.3.4 Penalized Regression Calibration (PRC)
The Penalized Regression Calibration (PRC) method (Signorelli et al.,, 2021) uses mixed-effects models to summarize the $Q$ longitudinal covariates into predicted random effects, which are then used as predictors of survival in a Cox model.
@@
where $\beta_{q}$ is a vector of population parameters shared across all subjects called fixed effects, $u_{qi}$ is a vector of subject-specific parameters called random effects that are assumed to follow a multivariate normal distribution, $\varepsilon_{qi}$ is a vector of Gaussian errors, and $W_{qi}(t_{ij})$ and $Z_{qi}(t_{ij})$ are design matrices containing the covariates relevant for $\beta_{q}$ and $u_{qi}$ respectively. Signorelli et al., (2021) also consider an alternative, more complex multivariate mixed model, which is not relevant for our model benchmarking. Specifically, in this article we consider a LMM with a random intercept and a random slope that depends on time $t_{ij}$:
@@
where $\beta_{q}=(\beta_{q0},\beta_{q1})$, $u_{qi}=(u_{qi0},u_{qi1})^{T}\sim N(0,\Sigma_{q})$, and $\varepsilon_{qi}\sim N(0,\sigma^{2}_{q}I)$. Notice that although Signorelli et al., (2021) and Signorelli, (2023) suggest to employ age at observation as time covariate in equation (8), here we choose to model $y_{qij}$ as a function of time on study $t_{ij}$ to make PRC directly comparable with MFPCCox, FunRSF and DynForest, as all of those methods model $y_{qij}$ on the time on study scale.
After model (7) has been estimated, a vector $\hat{u}_{qi}$ with the predicted random effects for covariate $q$ and subject $i$ is computed. In the case of model (8), that means that the predicted random intercept $\hat{u}_{qi0}$ and slope $\hat{u}_{qi1}$ are obtained as summaries of the vector $y_{qi}$. To predict survival, PRC uses a Cox model where the predicted random effects are used as predictors of survival together with the baseline covariates:
$h_{i}(t)=h_{0}(t)\exp\left\{\alpha x_{i}+\sum_{q=1}^{Q}\gamma_{q}\hat{u}_{qi0}+\sum_{q=1}^{Q}\theta_{q}\hat{u}_{qi1}\right\},\>i\in\mathcal{I}(\ell),$ (9)
where $\hat{u}_{qi0}$ and $\hat{u}_{qi1}$ denote the estimated random effects from model (8), and $\alpha$, $\gamma=(\gamma_{1},...,\gamma_{Q})$ and $\theta=(\theta_{1},...,\theta_{Q})$ are vectors of regression coefficients. Model (9) is estimated using penalized maximum likelihood.
The PRC method can be estimated using the R package pencal (Signorelli,, 2023). For PRC, the predicted survival for subject $i$ can be estimated using equation (3), where $\hat{\eta}_{i}$ denotes the estimated linear predictor for subject $i$ from model (9).
@@ 2.3.5 Functional Random Survival Forest (FunRSF)
The Functional Random Survival Forest (FunRSF) (Lin et al.,, 2021) approach uses MFPCA to summarize the trajectories described by the longitudinal covariates, and random survival forests to predict survival. The MFPCA modelling step is analogous to the one used by MFPCCox and described in equation (5). This step will yield a set of $K$ estimated MFPCA scores $\hat{\rho}_{ki},k=1,...,K$ for each subject $i\in\mathcal{I}(\ell)$, which are then used within a random survival forest (RSF, Iswaran et al.,, 2008) algorithm as time-independent covariates together with the baseline covariates $x_{i}$
The estimation of the RSF involves sampling $B$ bootstrap samples from a dataset that comprises information on the survival outcome, the baseline covariates and the MFPCA scores. For each boostrap sample, a random survival tree is estimated, using as splitting variables both the baseline covariates and the MFPCA scores. At each node of the tree, a fixed number $F$ of candidate splitting variables is selected among the union of the $P$ baseline covariates and $K$ MFPCA score variables. Each node is split by identifying the candidate variable whose splitting maximizes the log-rank test statistic, and the splitting is continued until a terminal node is reached. A node is terminal if further splitting would lead to child nodes that contain less than a fixed number $s$ of subjects (notice that both $F$ and $s$ are tuning parameters that can be determined by the user). Each terminal node $k$ in the $b$-th tree is then summarized by its cumulative hazard function (CHF) $\hat{H}_{bk}(t|\ell)$. The estimated CHF for subject $i$ is obtained by dropping $i$ through each tree to obtain its estimated CHF $\hat{H}_{b}(t|\ell,x_{i},\hat{\rho}_{i})$, and then averaging over the $B$ bootstrap samples to obtain the ”ensemble” CHF estimate
@@
which is then used to estimate $\hat{S}_{i}\left(t|\ell,\mathcal{H}_{i}(\ell)\right)=\exp\left\{\hat{H}(t|\ell,x_{i},\hat{\rho}_{i})\right\}$. MFPCCox can be estimated using the R code provided in Lin et al., (2021).
@@ 2.3.6 Dynamic Random Survival Forest (DynForest)
The Dynamic Random Survival Forest (DynForest, Devaux et al.,, 2022) method makes use of LMMs to summarize the $Q$ longitudinal covariates into predicted random effects that are used as predictors of survival within a RSF.
The estimation of the RSF starts with B bootstrap samples from a dataset that comprises information on the survival outcome, the baseline covariates and the longitudinal covariates. For each bootstrap sample, a survival tree is estimated as follows: at each node, a fixed number $F$ of candidate splitting variables is selected among the union of the $P$ baseline covariates and the $Q$ longitudinal covariates. For each longitudinal covariate, a linear mixed-effects model of the form given in equation (7) is estimated, and the longitudinal trajectory of covariate $q$ for subject $i$, $y_{qi}$, is then summarized by the predicted random effects in the same way as described in Section 2.3.4 for PRC. The predicted random effects from the candidate longitudinal covariates are then used as time-independent covariates alongside with the candidate baseline covariates as possible splitting variables for the given node. Similarly to FunRSF, the node is split choosing the covariate that maximizes the log-rank statistics. The construction of the tree continues until a terminal node that contains a pre-specified number of subjects $s$, and a prespecified number of events $e$ is reached. Lastly, the estimated CHF for subject $i$ is computed by dropping $i$ down each tree, obtaining their estimated CHF according to the $b$-th tree, $\hat{H}_{b}(t|\ell,x_{i},y_{i})$, and averaging those tree-specific CHFs to obtain the ensemble CHF estimate
$\hat{H}(t|\ell,x_{i},y_{i})=\frac{1}{B}\sum_{b=1}^{B}\hat{H}_{b}(t|\ell,x_{i},y_{i}),$ (11)
which is then used to estimate $\hat{S}_{i}\left(t|\ell,\mathcal{H}_{i}(\ell)\right)=\exp\left\{\hat{H}(t|\ell,x_{i},y_{i})\right\}$.
Notice that although the construction of the RSF in DynForest shares similarities with FunRSF, an important difference is that FunRSF uses a single MFPCA computation to summarize the longitudinal covariates into MFPCA scores prior to estimating the RSF, whereas DynForest estimates new LMMs within each tree and node (meaning that the same longitudinal covariate may be summarized by different predicted random effects across trees and nodes). In addition, DynForest can be used also in combination with competing risks, which we do not consider in this article as our goal is to compare the predictive performance of the aforementioned methods (most of which have not been developed to incorporate competing risks). TheDynForest model can be estimated using the R package DynForest (Devaux et al.,, 2023).

### v2

@@ 2.3.1 Static Cox model
$h_{i}(t)=h_{0}(t)\exp\left\{\alpha^{T}x_{i}+\gamma^{T}y_{i}(0)\right\},\>i\in\mathcal{I}(\ell),$ (2)
where $h_{i}(t)$ denotes the hazard for subject $i$ at time $t$, $h_{0}(t)$ is a non-parametric baseline hazard, and $\alpha$ and $\gamma$ are vectors of regression coefficients. The predicted conditional survival probability of equation (1) can be estimated by computing
@@
Admittedly, model (2) does not make use of the available longitudinal measurements to update predictions of survival: nevertheless, it is useful to evaluate the predictive performance of this model to quantify how accurate our predictions of survival would be if we were to ignore any information collected after baseline. By comparing the performance of this static prediction model to that of dynamic methods, which incorporate more and more repeated measurements as the landmark time increases, we can effectively quantify the extent to which adding longitudinal information to a RPM may improve predictive performance. Although the discrepancy between static and dynamic methods will be typically data-dependent, in general we can expect the difference between these two approaches to increase as the landmark time increases, as the information provided to model (2) becomes more and more outdated, while dynamic methods continue being fed more and more up to date patient information.
@@ 2.3.2 Landmarking
Landmarking (Van Houwelingen, , 2007) is a pragmatic approach to dynamic prediction that summarizes the trajectory of each longitudinal covariate up until the landmark time into a single summary measure per covariate. Two commonly-used summary measures are the last available observation (up to $\ell$) of the longitudinal covariate, an approach referred to as last observation carried forward (LOCF), and the average of all repeated measurements taken up until $\ell$. The summaries of the longitudinal covariates are then included as predictors in a Cox model alongside the baseline covariates.
Let $y_{qi}=(y_{qi1},y_{qi2},...,y_{qim_{i}})^{T}$ denote all repeated measurements of the $q$-th longitudinal covariate for subject $i\in\mathcal{I}(\ell)$, and denote by $y_{qi}^{\ell}$ the last non-missing value in $y_{qi}$ such that $t_{ik}\leq\ell$. The LOCF landmarking method proceeds to fit the following Cox model
$h_{i}(t)=h_{0}(t)\exp\left\{\alpha^{T}x_{i}+\sum_{q=1}^{Q}\gamma_{q}y_{qi}^{\ell}\right\},\>i\in\mathcal{I}(\ell),$ (4)
where $\alpha$ and $\gamma=(\gamma_{1},\gamma_{2},...,\gamma_{Q})^{T}$ are vectors of regression coefficients. The predicted survival for subject $i$ can be estimated using equation (3), where $\hat{\eta}_{i}$ is now the estimated linear predictor from model (4).
@@ 2.3.3 Multivariate Functional Principal Component Analysis Cox model (MFPCCox)
The Multivariate Functional Principal Component Analysis Cox (MFPCCox) model (Li and Luo, , 2019; Gomon et al., , 2024) uses Multivariate Functional Principal Component Analysis (MFPCA) to summarize the $Q$ longitudinal covariates into MFPCA scores, which are then used as predictors of survival in a Cox model.
@@
where $\psi_{k}(t)=\left(\psi_{k1}(t),...,\psi_{kQ}(t)\right)^{T}$ are $Q$-variate orthonormal eigenfunctions with associated MFPCA scores $\rho_{ki}$. The eigenfunctions are ordered by decreasing percentage of variance explained, and $K$ is chosen in such a way that the first $K$ eigenfunctions explain a certain proportion of the variance of the longitudinal covariates.
This is achieved by assuming that the observed $Y_{q}(t)$ is a noisy realization from an underlying random process $R_{q}(t)$ with mean $\mu_{q}(t)=E[R_{q}(t)]$ and covariance function $C_{q}(t,t^{\prime})=Cov(R_{q}(t),R_{q}(t^{\prime}))$: $Y_{q}(t)=R_{q}(t)+\epsilon_{q}(t)$, where $\epsilon_{q}(t)$ are i.i.d. gaussian errors. First, one needs to estimate the mean $\mu_{q}(t)$ and covariance $C_{q}(t,t^{\prime})$ functions using non-parametric estimators. Then, $C_{q}(t,t^{\prime})$ is decomposed to obtain the eigenfunctions, and the corresponding univariate FPCA scores are computed using the Principal component Analysis by Conditional Expectation (PACE) algorithm (Yao et al., , 2005). Lastly, estimates of the multivariate FPCA scores $\rho_{ki}$ and eigenfunctions $\psi_{kq}(t)$ in equation (5) are computed from the univariate FPCA scores (Happ and Greven, , 2018).
To predict survival, MFPCCox considers a Cox model where the estimated MFPCA scores are included in the linear predictor alongside the baseline covariates:
$h_{i}(t)=h_{0}(t)\exp\left\{\alpha^{T}x_{i}+\sum_{k=1}^{K}\gamma_{k}\hat{\rho}_{ki}\right\},\>i\in\mathcal{I}(\ell),$ (6)
where $\hat{\rho}_{ki}$ are the estimated MFPCA scores, and $\alpha$ and $\gamma=(\gamma_{1},\gamma_{2},...,\gamma_{K})^{T}$ are vectors of regression coefficients. The predicted survival for subject $i$ can be estimated using equation (3), where $\hat{\eta}_{i}$ is now the estimated linear predictor from model (6).
@@ 2.3.4 Penalized Regression Calibration (PRC)
The Penalized Regression Calibration (PRC) method (Signorelli et al., , 2021; Signorelli, , 2024) uses mixed-effects models to summarize the $Q$ longitudinal covariates into predicted random effects, which are then used as predictors of survival in a Cox model.
@@
where $\beta_{q}$ is a vector of population parameters shared across all subjects called fixed effects, $u_{qi}$ is a vector of subject-specific parameters called random effects that are assumed to follow a multivariate normal distribution, $\varepsilon_{qi}$ is a vector of Gaussian errors, and $W_{qi}(t_{ij})$ and $Z_{qi}(t_{ij})$ are design matrices containing the covariates relevant for $\beta_{q}$ and $u_{qi}$, respectively. Notice that Signorelli et al., (2021) also considered a more complex multivariate mixed model that is not relevant for the datasets considered in this paper.
Specifically, in this article we consider an LMM with a random intercept $u_{qi0}$ and a random slope $u_{qi1}$ associated to the time variable $t_{ij}$:
@@
where $\beta_{q}=(\beta_{q0},\beta_{q1})^{T}$, $u_{qi}=(u_{qi0},u_{qi1})^{T}\sim N(0,\Sigma_{q})$, and $\varepsilon_{qi}\sim N(0,\sigma^{2}_{q}I)$. Notice that although Signorelli et al., (2021) and Signorelli, (2024) suggest to employ age at observation as time covariate in equation (8), here we choose to model $y_{qij}$ as a function of time on study $t_{ij}$ to make PRC directly comparable with MFPCCox, FunRSF and DynForest, as all of those methods model $y_{qij}$ on the time on study scale.
After model (7) has been estimated, PRC proceeds to compute the vector with the predicted random effects $\hat{u}_{qi}$ for each covariate. As an example, for model (8) PRC computes a predicted random intercept $\hat{u}_{qi0}$ and slope $\hat{u}_{qi1}$, which are subsequently used as summaries of the vector $y_{qi}$.
To predict survival, PRC uses a Cox model where the predicted random effects are used as predictors of survival together with the baseline covariates:
$h_{i}(t)=h_{0}(t)\exp\left\{\alpha^{T}x_{i}+\sum_{q=1}^{Q}\gamma_{q}\hat{u}_{qi0}+\sum_{q=1}^{Q}\theta_{q}\hat{u}_{qi1}\right\},\>i\in\mathcal{I}(\ell),$ (9)
where $\hat{u}_{qi0}$ and $\hat{u}_{qi1}$ denote the estimated random effects from model (8), and $\alpha$, $\gamma=(\gamma_{1},...,\gamma_{Q})^{T}$ and $\theta=(\theta_{1},...,\theta_{Q})^{T}$ are vectors of regression coefficients. Model (9) is estimated using penalized maximum likelihood. The predicted survival for subject $i$ can be estimated using equation (3), where now $\hat{\eta}_{i}$ denotes the estimated linear predictor from model (9).
@@ 2.3.5 Functional Random Survival Forest (FunRSF)
The Functional Random Survival Forest (FunRSF) approach (Lin et al., , 2021) uses MFPCA to summarize the trajectories described by the longitudinal covariates, and a random survival forest (RSF, Iswaran et al., , 2008) to predict survival.
The MFPCA modelling step is analogous to the one used by MFPCCox and described in Section 2.3.3. This step yields $K$ estimated MFPCA scores $\hat{\rho}_{ki},k=1,...,K$ for each subject $i\in\mathcal{I}(\ell)$, which are then used within the RSF as time-independent covariates together with the baseline covariates $x_{i}$.
The estimation of the RSF involves sampling $B$ bootstrap samples from a dataset that comprises information on the survival outcome, the baseline covariates and the MFPCA scores. For each boostrap sample, a random survival tree is estimated using as splitting variables both the baseline covariates, and the MFPCA scores. At each node of the tree, a fixed number $F$ of candidate splitting variables is selected among the union of the $P$ baseline covariates and $K$ MFPCA scores. Each node is split by identifying the candidate variable whose splitting maximizes the log-rank test statistic, and the splitting is continued until a terminal node is reached. A node is terminal if further splitting would lead to child nodes that contain less than a fixed number $s$ of subjects. Each terminal node $k$ in the $b$-th tree is then summarized by its cumulative hazard function (CHF) $\hat{H}_{bk}(t|\ell)$. The estimated CHF for subject $i$ is obtained by dropping $i$ through each tree to obtain its estimated CHF $\hat{H}_{b}(t|\ell,x_{i},\hat{\rho}_{i})$, and then averaging over the $B$ bootstrap samples to obtain the ensemble CHF estimate
@@
which is then used to estimate $\hat{S}_{i}\left(t|\ell,\mathcal{H}_{i}(\ell)\right)=\exp\left\{\hat{H}(t|\ell,x_{i},\hat{\rho}_{i})\right\}$.
@@ 2.3.6 Dynamic Random Survival Forest (DynForest)
The Dynamic Random Survival Forest (DynForest) method (Devaux et al., , 2023, 2024) makes use of LMMs to summarize the $Q$ longitudinal covariates into predicted random effects that are used as predictors of survival within an RSF.
The estimation of the RSF starts by drawing B bootstrap samples from a dataset that comprises information on the survival outcome, the baseline covariates and the longitudinal covariates. For each bootstrap sample, a survival tree is estimated as follows: at each node, a fixed number $F$ of candidate splitting variables is selected among the union of the $P$ baseline covariates and the $Q$ longitudinal covariates. For each candidate longitudinal covariate, an LMM of the form given in equation (7) is estimated, and the longitudinal trajectory of covariate $q$ for subject $i$, $y_{qi}$, is then summarized by the predicted random effects $\hat{u}_{qi}$ in the same way as done by PRC (see Section 2.3.4). The predicted random effects from the candidate longitudinal covariates are then used as time-independent covariates alongside the candidate baseline covariates as possible splitting variables for the given node. Similarly to FunRSF, the node is split choosing the candidate covariate that maximizes the log-rank statistics. The construction of the tree continues until a terminal node that contains a pre-specified number of subjects $s$, and a prespecified number of events $e$ is reached. Lastly, the estimated CHF for subject $i$ is computed by dropping the individual down each tree, obtaining their estimated CHF according to the $b$-th tree, $\hat{H}_{b}(t|\ell,x_{i},\hat{u}_{i})$, and averaging those tree-specific CHFs to obtain the ensemble CHF estimate
$\hat{H}(t|\ell,x_{i},\hat{u}_{i})=\frac{1}{B}\sum_{b=1}^{B}\hat{H}_{b}(t|\ell,x_{i},\hat{u}_{i}),$ (11)
which is then used to estimate $\hat{S}_{i}\left(t|\ell,\mathcal{H}_{i}(\ell)\right)=\exp\left\{\hat{H}(t|\ell,x_{i},\hat{u}_{i})\right\}$.
Notice that although the construction of the RSF in DynForest shares similarities with FunRSF, an important difference is that FunRSF uses a single MFPCA computation to summarize the longitudinal covariates into MFPCA scores prior to estimating the RSF, whereas DynForest estimates new LMMs within each tree and node (meaning that the same longitudinal covariate may be summarized by different predicted random effects across trees and nodes). Differently from MFPCCox, PRC and FunRSF, DynForest can handle competing risks (which we do not address this in this article, since most multi-step methods have not yet been extended to the competing risks setting).
2.3.7 Statistical software
The static Cox model and LOCF landmarking model can be estimated using the coxph function from R package survival (Therneau and Grambsch, , 2000). Dedicated R packages are available for PRC and DynForest: PRC can be estimated using the R package pencal (Signorelli, , 2024), and DynForest using the R package DynForest (Devaux et al., , 2024). As concerns the remaining two methods, MFPCCox can be estimated using the R code provided in Li and Luo, (2019) and FunRSF using the R code provided in Lin et al., (2021).
All computations presented in this study were performed on an AMD EPYC 7662 processor with 2 GHz CPU.